In a calorimetry experiment, butan-1-ol was completely combusted to heat 150.0 mL of water from 20.0◦C to 70.0◦C. The molar heat of combustion of butan-1-ol is 2670 kJmol–1.

If only 65.0% of the heat released by the combustion was absorbed by the water, what volume of carbon dioxide gas was produced at 25◦C and 100 kPa?


1
Expert's answer
2021-09-09T02:44:01-0400

∆T = 70-20 = 50 + 273 = 323

∆H = 2670 kJmol-1

1 kpa = 0.00987 atmosphere = 100 × 0.00987 = 0.987atm

2670 = 0.987 × 323 × X

= 8.38L